Given f(x)= 4x-3 and g(x)= 2x^2-5 find f[g(-1)] and g[f(-1)]
i dont know how to do these. thanks.

Answer by fractalier(1829)

f[g(-1)] = f(2(-1)^2 - 5) = f(-3) = -15
g[f(-1)] = g(4(-1) - 3) = g(-7) = 2(-7)^2 - 5 = 93